Two railroad cars are being linked by striking upon each other. The first car has a mass of 150 tonne and an initial velocity of 1.08 kph due east, while the second has a mass of 110 tonne and a velocity of 0.432 kph due west. What is their final velocity in m/s? 


1
Expert's answer
2021-10-31T18:15:43-0400

Apply momentum conservation:


"m_1v_1+m_2v_2=(m_1+m_2)u,\\\space\\
u=\frac{m_1v_1+m_2v_2}{m_1+m_2}=\frac{150·1.08+110·(-0.432)}{150+110}=0.44\text{ kph}."

The direction is due east.